Soybean meal is 12% protein; cornmeal is 6% protein. How many pounds of each should be mixed together in order to get 240-lb mixture that is 10% protein?
How many pounds of the cornmeal should be in the mixture?
How many pounds of the soybean meal should be in the mixture?

Answer by jorel1380(3719)

240x.1=24
m+n=240
.12m+.06n=24
12m+6n=2400
6m+6n=1440
6m=960
m=160
n=80
You need 120 lbs. of soybean and 80 lbs. of cornmeal for your mix..